OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = -2x - 8 and contains the point (4, 4).

OpenStudy (yrelhan4):

whats the slope of this line?

OpenStudy (anonymous):

The one that it's perpendicular to is -2. I don't know the one that it's asking about. I know it's the opposite reciprocal, but I messed up on it. >.<

OpenStudy (yrelhan4):

its the opposite reciprocal with a negative sign.

OpenStudy (yrelhan4):

oh thats what it means. lol. where are you stuck?

OpenStudy (anonymous):

So, y = 2x + 4 ?

OpenStudy (anonymous):

I dont know :O lol I'm all confused.

OpenStudy (yrelhan4):

y=mx+c is the basic eq of a line, right?

OpenStudy (anonymous):

y = 1/2x + 2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Are you positive?

OpenStudy (yrelhan4):

so, if the given line has a slope -2. the line perpendicular to it has a slope 1/2. thats a result. you have to memorize it. m1*m2=-1 for 2 perpendicular lines.

OpenStudy (anonymous):

i think you need to understand the beauty that the slope(some geometric thing) can be represented by a single number!

OpenStudy (yrelhan4):

now, the line you need is y=1/2x+c. you dont know c. now thats the reason you are given point 4,4. plug them into y=1/2x+c, find c, and you have the line.

OpenStudy (anonymous):

Lol okay, thanks guys :P

OpenStudy (anonymous):

 try to understand the above picture.

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

it tells you exactly the equation m1*m2=-1 for 2 perpendicular lines

OpenStudy (anonymous):

you understand how slope is represented by a number right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

in my picture, m1 is m and m2 is -1/m. you can see those are perpendicular

OpenStudy (anonymous):

@yrelhan4 can you give him a medal that way you both get one?